# VR-Robo: A Real-to-Sim-to-Real Framework for Visual Robot Navigation and Locomotion

### v1

@@ Abstract
Figure 1: Our VR-Robo introduces a Real-to-Sim-to-Real framework. We generate photorealistic and physically interactive “digital twin” simulation environments for ego-centric visual navigation and locomotion. VR-Robo facilitates the RL policy training of legged agents with exploration capability in complex scenarios. Notably, our framework achieves RGB-only zero-shot sim-to-real policy transfer to the real world. Please refer to the supplementary video for simulation and real-world experiment results.
@@ I Introduction
Exploring, perceiving, and interacting with the physical real world is crucial for legged robotics, which supports many applications such as household service and industrial automation. However, conducting real-world experiments poses considerable challenges due to safety risks and efficiency constraints. Consequently, training in simulation [44, 31, 40, 33, 24, 52, 15] has emerged as an effective alternative, allowing robots to experience diverse environmental conditions, safely explore failure cases, and learn directly from their actions. Despite these advantages, transferring policies learned in simulated environments to real-world remains challenging, owing to the significant Sim-to-Real reality gap [25, 42, 44].
@@
Substantial efforts have been made to narrow the Sim-to-Real gap in legged locomotion. Previous studies have employed cross-domain depth images to train agents in simulation, achieving impressive zero-shot policy transfer for both quadruped [61, 9, 17, 50] and humanoid [62, 28] robots. To further integrate RGB perception into the Sim-to-Real pipeline, LucidSim [56] leverages generative models within simulation [44] for visual parkour. However, these depth-based and generation-based visual policies are predominantly constrained to low-level locomotion tasks, primarily because standard simulators struggle to replicate the visual fidelity and complex geometry of real-world environments.
@@
Recently, advanced neural reconstruction techniques such as Neural Radiance Fields (NeRF) [32] and 3D Gaussian Splatting (3DGS) [22] have emerged as promising solutions for creating Real-to-Sim “digital twins” from real-world data. Nonetheless, most existing approaches [3, 35, 29, 27, 36, 51, 54] focus on enhancing photorealism, offering limited support for physical interaction with complex terrains. Moreover, the simulations in these works often lack mechanisms for robust environment exploration, thus constraining their deployment in complex real-world scenarios that demand dynamic interaction.
@@
To address these challenges, we introduce VR-Robo, a novel Real-to-Sim-to-Real framework that enables realistic, interactive Real-to-Sim simulation and reinforcement learning (RL) policy training for legged robot navigation and locomotion, minimizing the Sim-to-Real gap. Given the multi-view images, we employ planar-based [6, 19, 7] 3DGS [22] and foundation-model priors [55] to reconstruct geometry-consistent scenes and objects. We then propose a GS-mesh hybrid representation with coordinate alignment to create a “digital twin” simulation environment in Isaac Sim, which supports ego-centric visual perception and mesh-based physical interactions. To enable robust policy training, we introduce an agent-object randomization and occlusion-aware scene composition strategy, as illustrated in Figure 6.
With the reconstructed simulation, we adopt a two-stage training strategy for the visual RL policy. In the first stage, we train a low-level locomotion policy that enables the robot to perceive and utilize terrain like slopes and stairs to reach the high platforms. However, this policy alone does not address high-level ego-centric navigation tasks with RGB inputs. In the second stage, we train a high-level policy to explore the environment, identify target goals, and plan feasible paths. The agent utilizes ego-centric visual observations rendered via GS, interacts with the mesh in Isaac Sim, and updates its policy using rewards, as depicted in Figure 5. Through our extensive experiments, the trained policy demonstrates sufficient robustness to be transferred zero-shot to real-world scenarios using only RGB observations.
@@
    A Real-to-Sim-to-Real framework for robot navigation and locomotion. We propose to reduce the Sim-to-Real gap by training RL policies within a realistic and interactive simulation environment reconstructed directly from real-world scenarios.
@@
    Zero-shot Sim-to-Real RL policy transfer. Through extensive experimental evaluations, we demonstrate that VR-Robo produces effective navigation and locomotion policies in complex, real-world scenarios using RGB-only observations.
Figure 2: Our Real-to-Sim-to-Real Framework. We first reconstruct the geometry-consistent scenarios from the captured images with foundation model constraints. Then we build a realistic and interactive simulation environment with GS-mesh hybrid representation and occlusion-aware composition and randomization for policy training. Finally, we zero-shot transfer the RL policy trained in simulation into the real robot for ego-centric navigation and visual locomotion.
@@ II-A Sim-to-Real Policy Transfer
Transferring reinforcement learning (RL) policies trained in simulation to the real world remains a major challenge due to substantial domain gaps. Traditional simulator-based methods, such as domain randomization [48, 34, 43, 53, 38, 46] and system identification [20, 41, 1, 5], aim to reduce the Sim-to-Real gap by aligning simulations with physical setups. Recent work [26, 49, 8, 57] proposes leveraging conditional generative models to augment visual observations for more robust agent training. LucidSim [56], for example, incorporates RGB color perception into the Sim-to-Real pipeline to learn low-level visual parkour by augmenting image background. However, these approaches remain constrained by conventional simulators, which fail to capture the full breadth of real-world physics and visual realism necessary for high-level policy training and real-world deployment.
@@ II-B Real-to-Sim Scene Transfer
Recently, advances in scene representation and reconstruction such as Neural Radiance Fields (NeRF) [32] and 3D Gaussian Splatting (3DGS) [22] have facilitated the creation of high-fidelity digital twins that closely replicate real-world environments for Real-to-Sim scene transfer. For instance, NeRF2Real [3] integrates NeRF into simulation for vision-based bipedal locomotion policy training, yet lacks physical interaction with reconstructed geometries. Meanwhile, RialTo [45] augments digital twins with articulated USD representations to enable manipulative interactions. More recent works employ 3DGS to generate realistic simulations [21] for both robot manipulation [29, 27, 36, 51] and navigation [35, 54]. In contrast, VR-Robo is designed to produce photorealistic and physically interactive “digital twin” environments specifically tailored for ego-centric visual locomotion learning.
@@ III Task definition
Figure 3: Task Definition. The robot needs to reach the target cone of a specified color, relying on RGB observations and proprioception.
@@
IV A Real-to-Sim-to-Real System for Ego-Centric Robot Locomotion
@@ IV-A Geometry-Consistent Reconstruction
Given multi-view RGB images of a scene with corresponding camera poses from COLMAP [39], we aim to generate a photorealistic and physically interactive simulation environment for agent policy training, thereby minimizing the sim-to-real gap. To this end, we first reconstruct a high-quality, geometry-aware 3D environment using planar-based 3D Gaussian Splatting (GS) [22] with geometric regularization.
Gaussian-based 3D Scene Representation. 3D Gaussian Splatting (3DGS) represents the scene as a set of Gaussian primitives. In particular, each Gaussian splat $\mathcal{G}_{i}(\mathbf{x})$ is parameterized by its mean $\mu_{i}\in\mathbb{R}^{3}$, a 3D covariance matrix $\Sigma_{i}\in\mathbb{R}^{3\times 3}$, an opacity term $o_{i}$, and a color term $c_{i}$, as follows:
@@
$\mathbf{c}(x)=\sum_{i\in N}T_{i}c_{i}\alpha_{i}(\mathbf{x}),\quad T_{i}=\prod_{i=1}^{i-1}\left(1-\alpha_{i}(\mathbf{x})\right).$ (2)
where $\alpha_{i}(\mathbf{x})=o_{i}\mathcal{G}_{i}(\mathbf{x})$ denotes the alpha of the Gaussian $\mathcal{G}_{i}$ at $\mathbf{x}\in\mathbb{R}^{3}$. Meanwhile, the view-dependent color $c_{i}$ of $\mathcal{G}_{i}$ is represented by its spherical harmonics $(\mathrm{SH})$ [23] coefficients.
Flattening 3D Gaussians for Geometric Modeling. Inspired by [19, 7], we flatten the 3D Gaussian ellipsoid with covariance $\Sigma_{i}=R_{i}S_{i}S_{i}^{T}R_{i}^{T}$ into 2D flat planes to enhance the scene geometry modeling by minimizing its shortest axis scale $S_{i}=$ $\operatorname{diag}\left(s_{1},s_{2},s_{3}\right)$:
$\mathcal{L}_{\mathtt{scale}}=\frac{1}{N}\sum_{i\in N}\left\|\min\left(s_{1},s_{2},s_{3}\right)\right\|,$ (3)
@@
$\bm{D}(p)_{\mathtt{render}}=\frac{\mathcal{D}}{\mathcal{N}(p)K^{-1}\tilde{p}}.$ (4)
@@
Foundation Model Geometric Prior Constraints. In textureless regions (e.g., ground and walls), photometric loss tend to be insufficient. To improve surface reconstruction in these areas, we employ an off-the-shelf monocular depth estimator [55] to provide dense depth prior. We address the inherent scale ambiguity between the estimated depths and the actual scene geometry by comparing them to sparse Structure-from-Motion (SfM) points, following [47, 10, 58]. Specifically, we align the scale of the monocular depth map $\bm{D}_{\mathtt{mono}}$ with the SfM points projected depth map $\bm{D}_{\mathtt{sfm}}$ by solving the per-image scale $\bm{s}$ and shift $\bm{t}$ parameters using linear regression:
$\hat{\bm{s}},\hat{\bm{t}}=\underset{\bm{s},\bm{t}}{\arg\min}\sum_{p\in\bm{D}_{\mathtt{sfm}}}\left\|\left(\bm{s}*\bm{D}(p)_{\mathtt{mono}}+\bm{t}\right)-\bm{D}(p)_{\mathtt{sfm}}\right\|_{2}^{2}.$ (5)
Finally, we use the re-scaled monocular depth map $\hat{\bm{s}}*\bm{D}(p)_{\mathtt{mono}}+\hat{\bm{t}}$ to regularize the rendered depth map $\bm{D}(p)_{\mathtt{render}}$. Similarly, we adopt an off-the-shelf monocular normal estimator [2] to provide dense normal regularization to the rendered normal map $\mathcal{N}_{\mathtt{render}}$ for accurate geometric modeling.
@@
$\mathcal{L}_{\mathtt{mv}}=\frac{1}{\|\mathcal{P}\|}\sum_{\mathbf{p}\in\mathcal{P}}\sum_{p\in\mathbf{p}}(1-NCC(\hat{\mathbf{I}}(\mathcal{H}p)_{\mathtt{render}},\mathbf{I}(p)_{\mathtt{render}})).$ (6)
@@ IV-B Building Realistic and Interactive Simulation
To enable the agent-environment interaction, we integrate a GS-mesh hybrid scene representation into the Isaac Sim with coordinate alignment. We further leverage agent-object randomization and occlusion-aware scene composition to advance robust visual policy training.
GS-mesh Hybrid Representation. To enhance the realism and interactivity of our simulation environment, we introduce a hybrid scene representation that combines our Gaussian-based model with its corresponding scene mesh primitives.
1) The Gaussian representation generates photorealistic visual observations from the robot’s ego-centric viewpoints. Because the robot camera differs from the camera used for 3D scene reconstruction, we first align the intrinsic parameters between Sim and Real by calibrating the robot camera’s focal length and distortion parameters. We then obtain the camera extrinsic within the simulation coordinate system by retrieving the ego-view position and quaternion-based orientation from Issac.
2) The Mesh representation facilitates physical interaction and precise collision detection. We render the depth for each input view and leverage a Truncated Signed Distance Function (TSDF) [11] fusion algorithm to build the corresponding TSDF field. We then extract the mesh from this TSDF field to enable physically accurate interactions within the simulation.
Coordinate Alignment. To align the reconstructed COLMAP coordinate system with the Isaac Sim environment, we first compute the homogeneous transformation matrix $T_{\mathtt{homo}}\in\mathbb{R}^{4\times 4}$ by manually matching four non-coplanar points as correspondence.
@@
$T_{\text{block}}=\arg\min_{T}\sum_{i=1}^{4}\|T\cdot\mathbf{p}_{i}^{\mathcal{B}}-\mathbf{p}_{i}^{\text{COLMAP}}\|^{2},$ (7)
@@
$T_{\text{cone}}=T_{\text{block}}\cdot\Delta\mathbf{p}_{\text{cone}}.$ (8)
Gaussian Attributes Adjustment. After applying the coordinate transformation, we must adjust the Gaussian attributes for accurate photometric rendering. Specifically, we first decompose $T_{\mathtt{homo}}$ into its rotation component $R_{\mathtt{homo}}\in\mathbb{R}^{3\times 3}$, translation vector $t_{\mathtt{homo}}\in\mathbb{R}^{3}$, and scale factor $s_{\mathtt{homo}}\in\mathbb{R}$. In addition, the 3D covariance matrix $\Sigma$ of the Gaussian points is parameterized by a scaling matrix $S\in\mathbb{R}^{3}$ and a rotation matrix $R\in\mathbb{R}^{3\times 3}$, such that $\Sigma=RSS^{T}R^{T}$. The mean $\mu$, scaling $S$ and rotation $R$ are adjusted as follows:
$\mu^{\prime}=R_{\mathtt{homo}}\mu+t_{\mathtt{homo}},$ (9)
$S^{\prime}=S+log(s_{\mathtt{homo}}),$ (10)
$R_{\mathtt{norm}}=\frac{R_{\mathtt{homo}}}{s_{\mathtt{homo}}}\quad\Sigma^{\prime}=R_{\mathtt{norm}}\Sigma R_{\mathtt{norm}}^{T},$ (11)
Figure 5: RL policy training in the reconstructed simulation environment. The agent leverages the ego-view GS photorealistic rendering as visual observations and interacts with the GS-extracted mesh in the Issac Sim environment. The agent receives the image feature $I_{a}$ from ViT [12] encoder, proprioception $P_{a}$ from simulator sensors, a task-specific RGB command $C_{a}$, and last action as input, using an asymmetric actor-critic LSTM structure to output the velocity command $V_{\text{raw\_cmd}}$ for low-level policy control.
@@
$\mathbf{C}^{(\ell)^{\prime}}=D_{\ell}(\alpha,-\beta,\gamma)\mathbf{C}^{(\ell)}.$ (12)
@@
Occlusion-Aware Composition and Randomization. To generate 3D object assets for environment randomization, we first capture high-quality multi-view images of real-world objects and reconstruct them using the aforementioned pipeline. With the resulting mesh and 3D Gaussians, we employ an interactive mesh editor to obtain both the 3D bounding box and its corresponding transformation matrix $T_{\mathtt{bbox}}\in\mathbb{R}^{4\times 4}$. Since the object mesh and the 3D Gaussians share the same COLMAP coordinate system, we use the mesh bounding box to crop the object Gaussians and merge them into the environment coordinate space according to the transformation below:
$T_{\mathtt{obj}}^{\mathtt{COLMAP-env}}=T_{\mathtt{sim}}^{\mathtt{COLMAP-env}}\cdot T_{\mathtt{COLMAP-obj}}^{\mathtt{sim}}\cdot T_{\mathtt{bbox}}.$ (13)
As shown in  Figure 6, the merged object Gaussians is synchronized with the object mesh in Issac and can be rendered jointly with the environment Gaussians via a volumetric alpha-blending [22] process. By incorporating z-buffering [4] on the opacity attributes of the Gaussian points, our method achieves occlusion-aware scene composition with accurate visibility. This randomization strategy substantially increases the diversity of the environment, thereby enabling more robust and scalable agent training.
Figure 6: Agent-object Randomization and Scene Composition. At the beginning of each episode, we randomly sample the positions for the robot and three cones in the Isaac Sim environment (upper row). We synchronously merge the agent and object Gaussians into the environment and compose for joint rendering (lower row). Both mesh and Gaussian rendering results shown here are in the Bird’s-Eye View (BEV) with the same camera pose.
IV-C RL in Reconstructed Simulation
Given the simulation environment generated in Section IV-B, the next stage of VR-Robo focuses on training a robust locomotion policy via reinforcement learning, as illustrated in Figure 5. This policy is designed to address diverse goal-tracking tasks under varying environmental conditions. Due to the limited onboard computational resources of real robots, we adopt a two-level asynchronous control strategy: the high-level policy operates at 5 Hz, while the low-level policy executes at 50 Hz. The high-level policy communicates with the low-level policy via velocity commands, represented as $V_{\text{cmd}}=(V_{x},V_{y},V_{yaw})$. All experiments are conducted primarily on the Unitree Go2 quadruped robot.
Low-Level Policy. The first step is to train the low-level policy, similar to previous locomotion works [50, 59]. The low-level policy takes velocity commands $V_{\text{cmd}}=(V_{x},V_{y},V_{\text{yaw}})$ along with robot proprioception as input, and outputs the desired joint positions. The actor network is composed of a lightweight LSTM layer combined with multiple MLP layers. Further details on the state and action spaces, reward design, network architecture, and terrain settings can be found in Appendix -C.

### v3

@@ Abstract
Index Terms:
Legged Robots, AI-Based Methods, Vision-Based Navigation.
@@ I Introduction
Exploring, perceiving, and interacting with the physical real world is crucial for legged robotics, which supports many applications such as household service and industrial automation. However, conducting real-world experiments poses considerable challenges due to safety risks and efficiency constraints. Consequently, training in simulation[1, 2, 3] has emerged as an effective alternative, allowing robots to experience diverse environmental conditions, safely explore failure cases, and learn directly from their actions. Despite these advantages, transferring policies learned in simulated environments to real world remains challenging, owing to the significant Sim-to-Real reality gap[4, 1].
@@
Substantial efforts have been made to narrow the Sim-to-Real gap in legged locomotion. Previous studies have employed cross-domain depth images to train agents in simulation, achieving impressive zero-shot policy transfer for both quadruped[5, 6, 7, 8] and humanoid[9] robots. To further integrate RGB perception into the Sim-to-Real pipeline, LucidSim[10] leverages generative models within simulation[1] for visual parkour. However, these depth-based and generation-based visual policies are predominantly constrained to low-level locomotion tasks, primarily because standard simulators struggle to replicate the visual fidelity and complex geometry of real-world environments.
@@
Recently, advanced neural reconstruction techniques such as Neural Radiance Fields (NeRF)[11] and 3D Gaussian Splatting (3DGS)[12] have emerged as promising solutions for creating Real-to-Sim “digital twins” from real-world data. Nonetheless, most existing approaches[13, 14, 15, 16, 17] focus on enhancing photorealism, offering limited support for physical interaction with complex terrains. Moreover, the simulations in these works often lack mechanisms for robust environment exploration, thus constraining their deployment in complex real-world scenarios that demand dynamic interaction.
Figure 1: Our VR-Robo introduces a unified real-to-sim-to-real framework. We generate photorealistic and physically interactive “digital twin” simulation environments for visual policy training, and enable zero-shot real-world deployment.
@@
To address these challenges, we introduce VR-Robo, a novel Real-to-Sim-to-Real framework that enables realistic, interactive Real-to-Sim simulation and reinforcement learning (RL) policy training for legged robot navigation and locomotion. Given the multi-view images, we employ 3DGS[12] and foundation-model priors[18] to reconstruct geometry-consistent scenes and objects. We then propose a GS-mesh hybrid representation with coordinate alignment to create a “digital twin” simulation environment in Isaac Sim, which supports ego-centric visual perception and mesh-based physical interactions. To enable robust policy training, we introduce an agent-object randomization and occlusion-aware scene composition strategy, as illustrated in Figure 7.
@@
    A Real-to-Sim-to-Real framework for robot navigation and locomotion. We propose to reduce the Sim-to-Real gap by training RL policies within a realistic and interactive simulation environment reconstructed only from RGB images captured from real-world.
@@
    Zero-shot Sim-to-Real RL policy transfer. Through extensive experimental evaluations, we demonstrate that VR-Robo produces effective navigation and locomotion policies in complex, real-world scenarios using RGB-only observations. In addition, we have some empirical findings for RL policy training, including the generalization ability, sample complexity, and architectural choices.
Figure 2: Diverse experiment settings including scenes, conditions, and robot types (better viewed when zoomed in).
@@ II-A Sim-to-Real Policy Transfer
Transferring reinforcement learning (RL) policies trained in simulation to the real world remains a major challenge due to substantial domain gaps. Traditional simulator-based methods, such as domain randomization[19, 20, 21] and system identification[22, 23], aim to reduce the Sim-to-Real gap by aligning simulations with physical setups. Recent work[24, 25] proposes leveraging conditional generative models to augment visual observations for more robust agent training. LucidSim[10], for example, incorporates RGB color perception into the Sim-to-Real pipeline to learn low-level visual parkour by generating and augmenting image background. However, these approaches remain constrained by conventional simulators, which fail to capture the full breadth of real-world physics and visual realism necessary for high-level policy training and real-world deployment.
@@ II-B Real-to-Sim Scene Transfer
Recently, advances in scene representation and reconstruction such as Neural Radiance Fields (NeRF)[11] and 3D Gaussian Splatting (3DGS)[12] have facilitated the creation of high-fidelity digital twins that closely replicate real-world environments for Real-to-Sim scene transfer. For instance, NeRF2Real[13] integrates NeRF and mesh into a simulation for vision-based bipedal locomotion policy training, but struggles to achieve real-time rendering, terrain climbing, and environment exploration, thus limiting their practical application. Meanwhile, RialTo[26] augments digital twins with articulated USD representations to enable manipulative interactions. More recent works employ 3DGS to generate realistic simulations for both robot manipulation[14, 15, 16] and navigation[27, 17]. In contrast, VR-Robo is designed to produce photorealistic and physically interactive “digital twin” environments specifically tailored for ego-centric visual locomotion learning.
@@ III Task definition
Our designed task requires both high-level understanding and navigation, as well as low-level motion control. As shown in Figure 3, the task is defined as reaching the target cone of a specified color within a limited time horizon. Each scene features a terrain in the center composed of a high table and a slope or stair connecting the ground to the platform.
Figure 3: Task definition. The robot aims to reach the target cone of a specified color in a specified spatial location, relying solely on ego-view RGB observation and proprioception.
Figure 4: VR-Robo real-to-sim-to-real framework. We build a realistic and interactive simulation environment with GS-mesh hybrid representation and occlusion-aware composition & randomization for policy training. Finally, we zero-shot transfer the RL policy trained in simulation into the real robot for ego-centric navigation and visual locomotion.
@@
IV A Real-to-Sim-to-Real System for Visual Locomotion
@@ IV-A Geometry-Consistent Reconstruction
Given multi-view RGB images of a scene with corresponding camera poses from COLMAP[28], we aim to generate a photorealistic and physically interactive simulation environment for agent policy training. We decoupledly reconstruct the environment and the object, and then compose them into a single Gaussian scene for policy training.
Gaussian-based 3D Scene Representation. 3D Gaussian Splatting (3DGS) represents the scene as a set of Gaussian primitives as follows:
@@
$\mathbf{c}(x)=\sum_{i\in N}T_{i}c_{i}\alpha_{i}(\mathbf{x}),\quad T_{i}=\prod_{i=1}^{i-1}\left(1-\alpha_{i}(\mathbf{x})\right).\vspace{-1mm}$ (2)
where $\alpha_{i}(\mathbf{x})=o_{i}\mathcal{G}_{i}(\mathbf{x})$ denotes the alpha of the Gaussian $\mathcal{G}_{i}$ at $\mathbf{x}\in\mathbb{R}^{3}$. Meanwhile, the view-dependent color $c_{i}$ of $\mathcal{G}_{i}$ is represented by its spherical harmonics $(\mathrm{SH})$ coefficients.
Flattening 3D Gaussians for Geometric Modeling. Inspired by[29], we flatten the 3D Gaussian ellipsoid with covariance $\Sigma_{i}=R_{i}S_{i}S_{i}^{T}R_{i}^{T}$ into 2D flat planes to enhance the scene geometry modeling by minimizing its shortest axis scale $S_{i}=$ $\operatorname{diag}\left(s_{1},s_{2},s_{3}\right)$:
$\mathcal{L}_{\mathtt{scale}}=\frac{1}{N}\sum_{i\in N}\left\|\min\left(s_{1},s_{2},s_{3}\right)\right\|,\vspace{-1mm}$ (3)
@@
$\boldsymbol{D}(p)_{\mathtt{render}}=\frac{\mathcal{D}}{\mathcal{N}(p)K^{-1}\tilde{p}}.\vspace{-1mm}$ (4)
@@
Foundation Model Geometric Prior Constraints. In texture-less regions (e.g., ground and walls), photometric loss tend to be insufficient. We employ an off-the-shelf monocular depth estimator[18] to provide dense depth prior. We address the inherent scale ambiguity between the estimated depths and the actual scene geometry by comparing them to sparse Structure-from-Motion (SfM) points, following[31]. Specifically, we align the scale of the monocular depth map $\boldsymbol{D}_{\mathtt{mono}}$ with the SfM points projected depth map $\boldsymbol{D}_{\mathtt{sfm}}$ using linear regression:
$\hat{\boldsymbol{s}},\hat{\boldsymbol{t}}=\underset{\boldsymbol{s},\boldsymbol{t}}{\arg\min}\sum_{p\in\boldsymbol{D}_{\mathtt{sfm}}}\left\|\left(\boldsymbol{s}*\boldsymbol{D}(p)_{\mathtt{mono}}+\boldsymbol{t}\right)-\boldsymbol{D}(p)_{\mathtt{sfm}}\right\|_{2}^{2}.\vspace{-1mm}$ (5)
Finally, we use the re-scaled depth map $\hat{\boldsymbol{s}}\times\boldsymbol{D}(p)_{\mathtt{mono}}+\hat{\boldsymbol{t}}$ to regularize the rendered depth map $\boldsymbol{D}(p)_{\mathtt{render}}$. Similarly, we adopt an off-the-shelf monocular normal estimator[32] to provide dense normal regularization to the rendered normal map $\mathcal{N}_{\mathtt{render}}$ for accurate geometric modeling.
@@
$\mathcal{L}_{\mathtt{mv}}=\frac{1}{\|\mathcal{P}\|}\sum_{\mathbf{p}\in\mathcal{P}}\sum_{p\in\mathbf{p}}(1-NCC(\hat{\mathbf{I}}(\mathcal{H}p),\mathbf{I}(p))).\vspace{-1mm}$ (6)
@@ IV-B Building Realistic and Interactive Simulation
To enable the agent-environment interaction, we integrate a GS-mesh hybrid scene representation into the Isaac Sim with coordinate alignment. We further leverage agent-object randomization and occlusion-aware scene composition to advance robust and generalizable visual policy training.
GS-mesh Hybrid Representation. 1) The Gaussian representation generates photorealistic visual observations from the robot’s ego-centric viewpoints. We first align the intrinsic parameters between Sim and Real by calibrating the robot camera’s focal length and distortion parameters. We then obtain the camera extrinsic within the simulation coordinate system by retrieving the ego-view position and quaternion-based orientation from Issac Sim. 2) The Mesh representation facilitates physical interaction and precise collision detection. We render the depth map for each input view and leverage a Truncated Signed Distance Function (TSDF)[35] fusion algorithm to build the corresponding TSDF field. We then extract the mesh from this TSDF field to enable physically accurate interactions within the simulation.
Coordinate Alignment. To align the reconstructed COLMAP coordinate system with the Isaac Sim environment, we first compute the homogeneous transformation matrix $T_{\mathtt{homo}}\in\mathbb{R}^{4\times 4}$ by manually matching four non-coplanar points.
@@
$T_{\text{block}}=\arg\min_{T}\sum_{i=1}^{4}\|T\cdot\mathbf{p}_{i}^{\mathcal{B}}-\mathbf{p}_{i}^{\text{COLMAP}}\|^{2},\vspace{-2mm}$ (7)
@@
$T_{\text{cone}}=T_{\text{block}}\cdot\Delta\mathbf{p}_{\text{cone}}.\vspace{-1mm}$ (8)
Gaussian Attributes Adjustment. We first decompose $T_{\mathtt{homo}}$ into its rotation component $R_{\mathtt{homo}}\in\mathbb{R}^{3\times 3}$, translation vector $t_{\mathtt{homo}}\in\mathbb{R}^{3}$, and scale factor $s_{\mathtt{homo}}\in\mathbb{R}$. In addition, the 3D covariance matrix $\Sigma$ of the Gaussian points is parameterized by a scaling matrix $S\in\mathbb{R}^{3}$ and a rotation matrix $R\in\mathbb{R}^{3\times 3}$, such that $\Sigma=RSS^{T}R^{T}$. The mean $\mu$, scaling $S$ and rotation $R$ are adjusted as follows:
$\mu^{\prime}=R_{\mathtt{homo}}\mu+t_{\mathtt{homo}},\vspace{-1mm}$ (9)
$S^{\prime}=S+log(s_{\mathtt{homo}}),\vspace{-1mm}$ (10)
$R_{\mathtt{norm}}=\frac{R_{\mathtt{homo}}}{s_{\mathtt{homo}}}\quad\Sigma^{\prime}=R_{\mathtt{norm}}\Sigma R_{\mathtt{norm}}^{T},\vspace{-1mm}$ (11)
Figure 6: RL policy training in the reconstructed simulation environment. The agent leverages the ego-view GS photorealistic rendering as visual observations and interacts with the GS-extracted mesh in the Issac Sim environment.
@@
$\mathbf{C}^{(\ell)^{\prime}}=D_{\ell}(\alpha,-\beta,\gamma)\mathbf{C}^{(\ell)}.\vspace{-2mm}$ (12)
@@
Occlusion-Aware Composition and Randomization. With the resulting mesh and 3D Gaussians, we employ an interactive mesh editor to obtain both the 3D bounding box and its corresponding transformation matrix $T_{\mathtt{bbox}}\in\mathbb{R}^{4\times 4}$. Since the object mesh and the 3D Gaussians share the same COLMAP coordinate system, we use the mesh bounding box to crop the object Gaussians and merge them into the environment coordinate space according to the transformation below:
$T_{\mathtt{obj}}^{\mathtt{COLMAP-env}}=T_{\mathtt{sim}}^{\mathtt{COLMAP-env}}\cdot T_{\mathtt{COLMAP-obj}}^{\mathtt{sim}}\cdot T_{\mathtt{bbox}}.\vspace{-1mm}$ (13)
As shown in Figure 7, the merged object Gaussians are synchronized with the object mesh in Issac and can be rendered jointly with the environment Gaussians via a volumetric alpha-blending [12] process. Since they are separately reconstructed, there are no holes in the merged 3DGS scenes. By incorporating z-buffering [36], it is impossible to see those objects that are farthest away from the ego-viewer and behind other objects. Therefore, our method can achieve occlusion-aware scene composition with accurate visibility. This randomization strategy substantially increases the environment’s diversity, enabling more robust and scalable agent training.
Figure 7: Agent-object Randomization and Scene Composition. At the beginning of each episode, we randomly sample the positions for the robot and three cones in the Isaac Sim environment (upper row). We synchronously merge the agent and object Gaussians into the environment and compose for joint rendering (lower row).
IV-C Reinforcement Learning in Reconstructed Simulation
The next stage of VR-Robo focuses on training a robust locomotion policy via reinforcement learning, as illustrated in Figure 6. This policy is designed to address diverse goal-tracking tasks under varying environmental conditions. Due to the limited onboard computational resources of real robots, we adopt a two-level asynchronous control strategy: the high-level policy operates at 5 Hz, while the low-level policy executes at 50 Hz. The high-level policy communicates with the low-level policy via velocity commands, represented as $V_{\text{cmd}}=(V_{x},V_{y},V_{yaw})$. Main experiments are conducted and demonstrated on the Unitree Go2 quadruped robot.
Low-Level Policy. The first step is to train the low-level policy, similar to previous locomotion works[8, 37]. The low-level policy takes velocity commands $V_{\text{cmd}}=(V_{x},V_{y},V_{\text{yaw}})$ along with robot proprioception as input, and outputs the desired joint positions. The actor network is composed of a lightweight LSTM layer combined with multiple MLP layers.